OpenStudy (anonymous):

Find the standard form of the equation of the parabola with a focus at (0, -9) and a directrix y = 9.

OpenStudy (anonymous):

use this.. (x - h)^2 = 4p(y - k)

OpenStudy (anonymous):

4p?

OpenStudy (campbell_st):

1st the directrix is above the focus... so the parabola is concave down. next the distance between the focus and directrix is twice the focal length... a  the focal length a = 9. the vertex is a units above the focus so the vertex is (0, -9 + 9) so the vertex of the parabola is at (0,0) so using the vertex form where (h, k) is the vertex and a is the focal length \[(x - h)^2 = 4a(y - k)\] substituting you get \[(x - 0)2 = 4 \times 9 (y - 0)\] or \[x^2 = 36y\] yo get it into standard from just make y the subject. hope it helps